A passenger trains speed is 60 mi/h, and a freight speed is 40 mi/h. The passenger train travels the same distance in 1.5 h less time than the freight. How long does each train take to make the trip?
Answer by stanbon(75887)

A passenger trains speed is 60 mi/h, and a freight speed is 40 mi/h. The passenger train travels the same distance in 1.5 h less time than the freight. How long does each train take to make the trip?
-----
Passenger train DATA:
rate = 60 mph ; time = x-1.5 hrs ; distance = rt = 60(x-1.5) = 60x-90 miles
-------------------
Freight train DATA:
rate = 40 mph ; time = x hrs ; distance = rt = 40x miles
-------------------
Equation:
distance = distance
40x = 60x-90
-20x = -90
x = 4.5 hrs (freight train time)
---
x-1.5 = 3 hrs (passesnger train time)
==========================
Cheers,
Stan H.
================